# Star Attention: Efficient LLM Inference over Long Sequences

### v1

@@ 3.3 Trade-off between accuracy and speed
Figure 5, illustrates the effect of varying block sizes during context encoding with sequence length of 128K tokens. Larger block sizes correlate to higher accuracy with Star Attention. This trend is consistent across all sequence lengths in our experiments.
From our experiments, we observe that setting the block size to approximately one-quarter of the total sequence length strikes an optimal trade-off between accuracy and speed. However, for sequence lengths exceeding 128K, as shown in Table 2, we fix the block size at 32K tokens to prioritize speedup, allowing for some acceptable accuracy degradation. Similarly, for larger models such as the Llama-3.1-70B-Instruct, we limit the block size to 16K tokens. The choice of block size is dependent on the user on how much accuracy can be traded for improved speed. As the block size increases, Star Attention’s performance approaches that of full global attention, providing users with flexibility in balancing computational efficiency with accuracy. Additional details regarding the experimental setup are provided in Appendix C.2.
Figure 5: Effect of block size on the accuracy of Star Attention on the RULER benchmark with block sizes ranging from 4K to 32K tokens for Llama-3.1-8B instruct model at sequence length of 128K. In each setting, the anchor block size matches the context block size. The results indicate that larger context block sizes are positively correlated with improved accuracy.
Model Seq. Len. Block Size Ring-Attn Star-Attn
(K) (K) Acc. (%) $\Delta$ Acc. $\Delta$ Speedup
128 32 77.39 +1.23% 2.7x
256 32 74.44 -1.04% 10.8x
512 32 69.30 -9.71% 16.2x
Llama3-8B-Instruct, 1048K Gradient.ai (2024) 1024 32 63.70 -8.36% 16.9x
64 16 88.54 -1.63% 4.7x
Llama-3.1-70B-Instruct, 128K Meta-AI (2024) 128 16 65.29 -11.44% 8.7x
Table 2: Accuracy versus speed trade-off for Star Attention compared to Ring (Global) Attention on RULER. The $\Delta$ for star attention shows the relative accuracy degradation and the relative speedup compared to global attention. When the block size remains fixed and the as sequence length increases, Star Attention achieves exponential speedup over Ring (Global) Attention at the cost of slightly more accuracy degradation. It is upto the user to decided how much accuracy they want to trade-off for speed by setting the block size.
@@ 3.4 In-Depth Analysis on RULER Task Categories
Figure 6: Accuracy (%) of Star Attention using the Llama-3.1-8B-Instruct model on the 5 categories of tasks in RULER on sequence lengths of 32K. In all experiments, the block size and anchor block size are set to one-quarter of the total sequence length. For the NIAH and QA tasks, Star Attention retains upto 95-100% accuracy of the baseline. The Multi-Hop Tracing task becomes quite challenging since it requires inter-block communication. In aggregation tasks, Star Attention show significant improvement as distributed local attention helps the model in such summarization tasks. The trend is consistent to other sequence lengths as shown in Appendix E (Figure 8)
In this section we investigate the strengths and limitations of Star Attention, using different categories of tasks within RULER. The benchmark has five primary categories: Single-NIAH, Multi-NIAH, Multi-Hop Tracing, Aggregation, and Question Answering (QA). Figure 6 presents categorical results of RULER for the Llama-3.1-8B-Instruct model on sequence length of 32K. The trend is consistent across all sequence lengths (16K, 32K, 64K, and 128K), as detailed in Appendix E (Figure 8). Notably, Star Attention achieves scores nearly identical to global attention in Single-NIAH tasks. However, in more complex tasks such as Multi-NIAH and QA, it shows slight decline in performance, with reductions ranging from 1.6% to 4.9% in Multi-NIAH and 0.9% to 6.8% in QA tasks. Despite these challenges, Star Attention consistently retains overall 95-100% accuracy of global attention.
Tasks such as Multi-Hop Tracing and Aggregation necessitate an in-depth comprehension of context. Multi-Hop Tracing poses a significant challenge for Star Attention, as it requires the model to propagate information across multiple hops within the sequence, demanding effective inter-block communication. However, Star Attention lacks inter-block communication, relying solely on global attention between the query and segregated KV caches within each block. Due to this, the performance degradation is considerable compared to global attention.
Aggregation tasks, encompassing Common and Frequent Words Extraction, assess models’ ability to aggregate relevant information within long-range contexts. Star Attention yields significant performance improvements across all sequence lengths. This enhancement stems from Star Attention’s chunk-based context division, enabling local attention within each chunk to strengthen summarization capabilities. Effective chunk-wise summarization in Phase 1 facilitates global attention’s information aggregation in Phase 2.
@@ 4.1 Position and Content of Anchor Block
In this section, we explore the role of anchor blocks during Phase 1 that enables Star Attention to approximate global attention behavior. As outlined in Section 2.1, anchor blocks are crucial in managing the attention spikes generated at the start of each context block, helping Star Attention approximate global attention (see Table 3 ) Drawing from the hypotheses on sink tokens in Xiao et al. (2024), we consider two potential explanations for the effectiveness of anchor blocks: (1) the model may develop a bias toward the absolute position of the anchor block, or (2) the semantic content of the anchor block is essential for maintaining performance. To better understand how anchor blocks enable Star Attention to approximate global attention distribution, we test both the hypotheses. We conduct experiments on the Llama-3.1-8B-Instruct model, varying both the position and content of the anchor block. We evaluate two configurations: a block size of 16K for sequences of length 64K, and a block size of 32K for sequences of length 128K, in both the cases, with anchor block size matching the context block size.
@@
Experiments RULER-NIAH (%)
64K $\Delta$64k 128k $\Delta$128k
Global attention 99.50 - 98.49 -
No anchor block 60.11 -39.59% 73.75 -25.12%
Content set to first-block, position IDs are:
   randomly sampled from [0, current_block) 96.79 -2.72% 97.16 -1.35%
   same as previous block 97.35 -2.16% 96.80 -1.71%
   same as first block 97.61 -1.90% 97.54 -0.96%
Position IDs set to first-block, content is:
   constant token (ex: ‘ ’ or ‘ the’ or ‘.’ ) 0.00 -100.00% 0 -100.00%
   random tokens 90.55 -8.99% 82.63 -10.15%
   shuffled first block tokens 92.96 -6.57% 90.76 -3.26%
   first block tokens 97.61 -1.90% 94.94 -0.96%
Previous-block used as anchor 94.20 -5.33% 96.13 -2.40%
Table 3: Experiments on analyzing the impact of varying the position and content of the anchor block with the LLaMA-3.1-8B-Instruct model, with a block size of 16K for 64K sequence length, and 32K for 128K sequence lengths. In each setting, the size of the anchor block matches the context block size. The $\Delta$ for star attention shows the relative accuracy degradation compared to global attention. The experiments are categorized into 4 groups: (i) absence of anchor block; (ii) varying the position IDs; (iii) varying the content; (iv) varying both the position and the content. Results indicate that while the anchor block’s position is not critical, its content is essential for optimal performance.
Figure 7: Effect of anchor block size on the accuracy of Star Attention on the RULER-NIAH benchmark with the Llama-3.1 8B Instruct model. In these experiments, the context block size is fixed to 32K for sequence length 128K, respectively. Results indicate that larger anchor block sizes lead to improved accuracy. The observed trend holds for all sequence lengths in our experiments.
@@ 4.2 Size of Anchor block
As discussed in Section 3.3, larger block sizes improve the accuracy of Star Attention. In this section, we analyze the impact of varying anchor block size while maintaining a fixed block size of 32K for a sequence length of 128K. As illustrated in Figure 7, increasing the anchor block size enhances model accuracy, with the best performance observed when the anchor block size equals the context block size. Although Figure 3(b) demonstrates that attention spikes predominantly occur in the first few tokens, reducing the number of tokens in the anchor block leads to a substantial drop in performance. This suggests that a larger anchor block is critical for maintaining model accuracy, despite attention spikes being concentrated at the beginning of the sequence. This observation implies that the anchor block’s effectiveness is not solely due to its role in managing attention sinks but may involve other underlying factors. These findings remain consistent across both base and instruct models, as well as for all sequence lengths. Further investigation into why the anchor block size must be equivalent to the context block size is left for future work.
@@ 5 Conclusion
In this paper, we introduced Star Attention, a novel block-sparse attention mechanism designed to enable efficient inference on long sequences in transformer-based LLMs. The method operates in two phases: (1) context tokens are processed using blockwise-local attention, with the context segmented into blocks where each block is prefixed with an anchor block; and (2) then the query and response tokens attend to all prior cached tokens through sequence-global attention. Star Attention delivers up to 11x speedup over Ring Attention while maintaining 95-100% accuracy, significantly enhancing both memory efficiency and inference speed. Scaling Star Attention to longer sequences (up to 1M) and larger models, we observe even greater speedups while preserving similar levels of accuracy. Despite these advances, several open questions remain. The role and optimal size of anchor blocks relative to context blocks require further exploration. Additionally, while Star Attention performs effectively with block sizes set to one-quarter of the sequence length, accuracy degrades when using smaller blocks on longer sequences. Future work will focus on refining the anchor block mechanism and improving performance on more complex long-context tasks to enhance the scalability and robustness of Star Attention.
@@ Appendix A Star Attention Pseudo-code
Appendix B Evaluation Benchmarks
We evaluate Star Attention on the RULER benchmark which comprises 13 tasks covering domains such as Needle-in-a-Haystack (Retrieval), Multi-Hop Tracing, Aggregation, and Question Answering. Each task comprises 500 samples. For the ablations, we choose four Needle-In-A-Haystack (NIAH) tasks where Paul Graham essays serve as the distractor text (haystack): Single 2, Single 3, MultiKey 1, and MultiQuery. In these tasks, a key-value pair is concealed within a long context, and the model must identify the value corresponding to the key based on the provided input query. Table 4 presents the configurations of all the tasks in RULER.
@@ Appendix B Evaluation Benchmarks
Additionally, we also evaluate our method on the BABILong benchmark. In BABILong, we choose 5 tasks (shown in Table 5), each containing a 1000 samples. These tasks are generated by simulating a set of characters and objects engaged in various movements and interactions across multiple locations. Each interaction is represented by a factual statement, and the objective is to answer questions based on the facts derived from the current simulation.

### v2

@@ 3.4 Trade-off between accuracy and speed
Figure 5(a) illustrates the effect of varying block size during context encoding, with the sequence length fixed at 128K tokens. Larger block sizes lead to improved accuracy, highlighting the benefits of increased receptive fields for long-context comprehension.
Empirically, setting the block size to approximately one-quarter of the total sequence length strikes an effective trade-off between accuracy and speed. For sequence lengths exceeding 128K, we fix the block size at 32K tokens to prioritize inference speed. As shown in Figure 6(a), this configuration allows Star Attention to achieve substantial speedups over Ring Attention while incurring only modest accuracy degradation. For instance, on the RULER benchmark with Llama-3-8B-Instruct-1048K, Star Attention achieves up to 11$\times$ speedup while retaining accuracy comparable to Ring Attention. At 1M tokens, the speedup increases to 16.9$\times$ with an accuracy drop of just 5.32%.
These findings demonstrate that Star Attention offers flexible control over the accuracy-efficiency trade-off. Larger block sizes allow performance to approach that of global attention, while smaller blocks enable higher throughput for latency-sensitive applications. The appropriate configuration can thus be tuned based on available resources and task requirements. Additional experimental details are provided in Appendix B.
@@ 3.5 In-Depth Analysis on RULER Task Categories
To better understand the strengths and limitations of Star Attention, we analyze its performance across different task categories within the RULER benchmark. RULER comprises five categories: Single-NIAH, Multi-NIAH, Multi-Hop Tracing, Aggregation, and Question Answering (QA). Figure 6(b) reports category-wise accuracy using the Llama-3.1-8B-Instruct model at a sequence length of 32K and a block size of 8K. We observe consistent trends across all sequence lengths, as detailed in Appendix D.
Star Attention performs comparably to global attention in the Single-NIAH, Multi-NIAH, and QA categories. These tasks typically involve localized retrieval or reasoning, where attention primarily operates within or near a single context block. In contrast, Multi-Hop Tracing presents a greater challenge. It requires propagating information across multiple hops within the sequence, demanding effective inter-block communication. Since Star Attention restricts KV-cache access to the local block during context encoding, the model lacks a mechanism for long-range token-to-token aggregation in this phase. Consequently, performance degrades relative to global attention.
Interestingly, Star Attention shows substantial gains in Aggregation tasks, especially those involving frequency analysis or summarization over distributed spans. Its chunk-wise encoding facilitates local aggregation within blocks, which is later synthesized during the global query phase. This two-phase process proves advantageous in capturing common patterns without needing full global context at once. This analysis suggests that Star Attention is especially well-suited for retrieval and aggregation tasks, while highlighting opportunities for future work on cross-block communication.
(a) Accuracy vs. Speed Trade-off with Fixed 32K Block Size
(b) Accuracy across different RULER task categories at 32K sequence length
Figure 6: Performance of Star Attention compared to Global (Ring) Attention on RULER. (a) Accuracy vs speed trade-off as sequence length increases from 128K to 1M with block size fixed at 32K. Star Attention achieves up to 16.9$\times$ speedup with modest accuracy degradation. (b) Accuracy across five RULER task categories at 32K sequence length with 8K block size. Star Attention matches or improves upon the baseline in most tasks, with significant gains in aggregation.
@@ 4 Ablation Study
Experiments RULER-NIAH (%)
64K $\Delta$64k 128k $\Delta$128k
Global attention 99.50 - 98.49 -
No anchor block 60.11 -39.39% 73.75 -24.74%
Content set to first-block, position IDs are:
   randomly sampled from [0, current_block) 96.79 -2.71% 97.16 -1.33%
   same as previous block 97.35 -2.15% 96.80 -1.69%
   same as first block 97.61 -1.89% 97.54 -0.95%
Position IDs set to first-block, content is:
   constant token (ex: ‘ ’ or ‘ the’ or ‘.’ ) 0.00 -99.5% 0 -98.49%
   random tokens 90.55 -8.95% 82.63 -15.86%
   shuffled first block tokens 92.96 -6.54% 90.76 -7.73%
   first block tokens 97.61 -1.89% 94.94 -3.55%
Previous-block used as anchor 94.20 -5.3% 96.13 -2.66%
Table 4: Impact of anchor block position and content on Star Attention accuracy using Llama-3.1-8B-Instruct on RULER-NIAH at 64K and 128K sequence lengths. Each configuration has the anchor block size equal to the context block. The $\Delta$ values indicate absolute accuracy degradation relative to global attention. Results show that anchor content is critical, while position IDs haveminor effect. Missing or poorly constructed anchors lead to significant degradation.
@@ 4.1 Position and Content of Anchor Block
In this section, we explore the role of anchor blocks during Phase 1 that enables Star Attention to approximate global attention behavior. As outlined in Section 2.1, anchor blocks are crucial in managing the attention spikes generated at the start of each context block, helping Star Attention approximate global attention (see Table 4 ) Drawing from the hypotheses on sink tokens in Xiao et al. (2024b), we consider two potential explanations for the effectiveness of anchor blocks: (1) the model may develop a bias toward the absolute position of the anchor block, or (2) the semantic content of the anchor block is essential for maintaining performance. To better understand how anchor blocks enable Star Attention to approximate global attention distribution, we test both the hypotheses. We conduct experiments on the Llama-3.1-8B-Instruct model, varying both the position and content of the anchor block. We evaluate two configurations: a block size of 16K for sequences of length 64K, and a block size of 32K for sequences of length 128K, in both the cases, with anchor block size matching the context block size.
@@ 4.2 Size of Anchor block
As discussed in Section 3.4, larger block sizes improve the accuracy of Star Attention. In this section, we analyze the impact of varying anchor block size while maintaining a fixed block size of 32K for a sequence length of 128K. As illustrated in Figure 5(b), increasing the anchor block size enhances model accuracy, with the best performance observed when the anchor block size equals the context block size. Although Figure 3(b) demonstrates that attention spikes predominantly occur in the first few tokens, reducing the number of tokens in the anchor block leads to a substantial drop in performance. This suggests that a larger anchor block is critical for maintaining model accuracy, despite attention spikes being concentrated at the beginning of the sequence. This observation implies that the anchor block’s effectiveness is not solely due to its role in managing attention sinks but may involve other underlying factors. These findings remain consistent across both base and instruct models, as well as for all sequence lengths. Further investigation into why the anchor block size must be equivalent to the context block size is left for future work.
5 Related Work
To address the computational challenges of long-context inference in LLMs, various techniques have emerged to mitigate memory usage and enhance inference speed.
Blockwise and Distributed Attention Computation: Flash Attention (Dao et al., 2022; Dao, 2024) introduces a blockwise GPU-efficient implementation of exact attention, reducing both memory footprint and runtime. Building on this, distributed approaches such as Liu et al. (2024a) and Shyam et al. (2024) partition the computation of self-attention and feed-forward networks across multiple devices, employing sophisticated communication-computation overlap to improve scalability. General distributed strategies (Shoeybi et al., 2019; Huang et al., 2019; Li et al., 2023; Meta-AI, 2021) provide frameworks for dividing the computational load effectively across multiple accelerators. These methods, however, still compute dense global attention, which becomes prohibitively expensive at longer sequence lengths. Star Attention leverages the distributed nature of these approaches but reduces attention complexity through a two-phase block-sparse approximation that avoids computing the full attention matrix.
Sparse Attention: Sparse attention methods reduce the quadratic complexity of self-attention through structured or learned sparsity patterns (Zhang et al., 2023; Tang et al., 2024; Child et al., 2019), achieving linear or log-linear scaling in sequence length (Dai et al., 2019; Qin et al., 2024). Beltagy et al. (2020) introduced sliding window attention combined with global tokens, which was adapted by Xiao et al. (2024b) for streaming generation via attention sinks. Jiang et al. (2024) focuses on identifying and leveraging dynamic sparse patterns, particularly to accelerate the pre-filling stage. More recently, Titans (Behrouz et al., 2024) augment LLMs with neural memory modules for long-horizon reasoning. Star Attention’s first phase is conceptually similar to streaming methods, but differs by utilizing global attention during decoding—preserving compatibility with pretrained models without retraining.
Memory Optimization: Maintaining the KV cache during autoregressive decoding is a major memory bottleneck. KV cache compression (Ge et al., 2024; Munkhdalai et al., 2024; Sun et al., 2024; Liu et al., 2024b; Wu et al., 2024) and low-rank approximation methods (Srebro & Jaakkola, 2003) have been proposed to trade precision for reduced memory. Recent systems explore eviction-based memory management strategies that allow LLMs to operate over virtually infinite contexts (Zhao et al., 2024; Han et al., 2024; Xiao et al., 2024a), often requiring architecture changes or specialized runtime support. Star Attention is orthogonal to these methods and can be integrated with them to further enhance inference efficiency.
@@ 6 Conclusion
In this paper, we introduced Star Attention, a novel block-sparse attention mechanism designed to enable efficient inference on long sequences in transformer-based LLMs. The method operates in two phases: (1) context tokens are processed using blockwise-local attention, with the context segmented into blocks where each block is prefixed with an anchor block; and (2) then the query and response tokens attend to all prior cached tokens through sequence-global attention. Star Attention delivers up to 11x speedup over Ring Attention while maintaining 97-100% accuracy, significantly enhancing both memory efficiency and inference speed. Despite these advances, several open questions remain. The role and optimal size of anchor blocks relative to context blocks require further exploration. Additionally, while Star Attention performs effectively with block sizes set to one-quarter of the sequence length, accuracy degrades when using smaller blocks on longer sequences. Future work will focus on refining the anchor block mechanism and improving performance on more complex long-context tasks to enhance the scalability and robustness of Star Attention.
@@ Appendix A Star Attention Pseudo-code
Appendix B Experiment Details
B.1 Baseline Comparison
Model Seq. Len. Block Size Ring-Attn Star-Attn
(K) (K) Acc. (%) $\Delta$ Acc. $\Delta$ Speedup
128 32 77.39 +0.96% 2.7x
256 32 74.44 -0.77% 10.8x
512 32 69.30 -6.73% 16.2x
Llama3-8B-Instruct, 1048K Gradient.ai (2024) 1024 32 63.70 -5.32% 16.9x
64 16 88.54 -1.44% 4.7x
Llama-3.1-70B-Instruct, 128K Meta-AI (2024) 128 16 65.29 -7.47% 8.7x
Table 5: Accuracy versus speed trade-off for Star Attention compared to Ring (Global) Attention on RULER. The $\Delta$ for star attention shows the absolute accuracy degradation and the relative speedup compared to global attention. When the block size remains fixed and the as sequence length increases, Star Attention achieves exponential speedup over Ring (Global) Attention at the cost of slightly more accuracy degradation. It is upto the user to decided how much accuracy they want to trade-off for speed by setting the block size.
Table 5 shows speedup obtained by Star Attention over the baseline on sequences over 128K tokens. For such long sequences, we freeze the block size to 32K sequences to optimize for speed. This setting shows upto 16.9x inference speedup with just 5.32% accuracy degradation compared to the baseline. Table 6 presents the time per sample for vanilla autoregressive generation, Ring Attention, and Star Attention across sequence lengths ranging from 16K to 128K. The results indicate that both Ring and Star Attention can process sequences up to 128K tokens on 8 A100 GPUs, whereas vanilla autoregressive inference encounters OOM issues beyond 64K tokens. For sequence lengths below 32K, vanilla inference is faster than the distributed attention mechanisms, primarily due to the GPU communication overhead incurred in the distributed setups. However, in long context scenarios i.e. on sequence lengths exceeding 32K tokens, Star Attention begins to demonstrate clear performance advantages. As demonstrated in Table 5, the speedup achieved by Star Attention increases significantly with longer sequence lengths.
Seq. Length Time Per Sample (s)
(K) Vanilla Ring Star
16 7 10 9
32 10 12 10
64 18 22 12
128 OOM 53 20
Table 6: Time per sample (seconds) for Llama3.1-8B-Instruct model with dense, ring, and star attention, using 8 A100 GPUs. Vanilla autoregressive generation encounters out-of-memory (OOM) at 128K sequence length. It performs best in short context scenarios (i.e. sequences upto 32K tokens) but in long context scenarios, star attention demonstrates significant speedup.
B.2 Hardware for Inference Speed
Model Size Seq. Length # GPUs # Workers
8B 16K - 128K 8 4
256K - 512K 16 8
1M 32 16
70B 16K - 32K 8 4
64K 16 4
128K 32 8
Table 7: Resources used for the speedup experiments
B.3 Prompt Templates
Prompt structure for base models:
1{context}{query}{answer_prefix}
Prompt structure for Llama-3 and Llama-3.1 Instruct models:
2
4
6
7{answer_prefix}
Appendix C Evaluation Benchmarks
RULER: This benchmark comprises 13 tasks covering domains such as Needle-in-a-Haystack (Retrieval), Multi-Hop Tracing, Aggregation, and Question Answering. Each task comprises 500 samples. For the ablations, we choose four Needle-In-A-Haystack (NIAH) tasks where Paul Graham essays serve as the distractor text (haystack): Single 2, Single 3, MultiKey 1, and MultiQuery. In these tasks, a key-value pair is concealed within a long context, and the model must identify the value corresponding to the key based on the provided input query. Table 8 presents the configurations of all the tasks in RULER.
@@ Appendix C Evaluation Benchmarks
BABILong: In BABILong, we choose 5 tasks (shown in Table 9), each containing a 1000 samples. These tasks are generated by simulating a set of characters and objects engaged in various movements and interactions across multiple locations. Each interaction is represented by a factual statement, and the objective is to answer questions based on the facts derived from the current simulation.